OpenStudy (anonymous):

what is the distance between -4 and 6

OpenStudy (ipwnbunnies):

6 is biggest than -4. Subtract -4 from 6.

OpenStudy (anonymous):

the distance between -4 and 6 is: \(\large 6-(-4)=6+4=10\)